Can someone help me on this problem?
I need to solve: 8x - 5 + 9x = 17x - 5
Thank you,
Lisa

Answer by funmath(2933)

Hi Lisa,
Solve: 8x - 5 + 9x = 17x - 5
8x+9x-5=17x-5
17x-5=17x-5
See how both sides match exactly?
This is an identity, the answer is all real numbers, because no matter what we put in for x both sides will always equal.
Happy Calculating!!!